OpenStudy (anonymous):

Find the volume V obtained by rotating the region bounded by the given curves about the specified axis.y = sin^2(x), y = 0, 0 ≤ x ≤ π; about the x−axis

OpenStudy (anonymous):

have you graphed the function yet? some people can just look at a problem and know which volume technique to use, but I still need to see a graph. It looks to me like we could use the disk method. What do you think?

OpenStudy (anonymous):

yes , it think it is a disk

OpenStudy (anonymous):

ok. So, if we call "y" f(x), then I think we can say that f(x) is always greater than y=0 over that interval. If you agree, then I think that makes it easier because we then don't have to worry about subtracting a g(x).

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi}\pi(f(x)^2)dx\]

OpenStudy (anonymous):

that is how I'd set it up